# Stability of Non-Hyperbolic Sets

### v1

@@ 5.3 Global curves of secondary tangencies
Let $(t^{*},a^{*})\in\mathcal{B}_{n}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. By Lemma 13 there exist $\epsilon>0$ and a smooth curve
$b_{n,n_{0}}:[t^{*}-\epsilon,t^{*}+\epsilon]\mapsto\mathbb{R}$
such that, for all $t\in[t^{*}-\epsilon,t^{*}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$. We want now to prove that this curve does not stop. Suppose it does.
Consider a sequence $t_{n}\in[t^{*}-\epsilon,t^{*}+\epsilon]$ of secondary tangencies of type $n_{0}$ that converges to a boundary point of $t^{*}-\epsilon,t^{*}+\epsilon]$. Since the set of tangencies is a close set, the limit point is again a tangency. Because of Lemma 14 one cannot have more than one limit point, indeed one can have only one tangency point while moving in the $a$ direction. By lemmas 15 and 16 this unique limit point cannot be on the boundary lines that define $\mathcal{B}_{n}$. Such a limit point is then contained in $\mathcal{B}_{n}$. One can then apply again Proposition 13 to this point and build a smooth continuous curve of maps with secondary tangencies of type $n_{0}$. This new courve $b^{\prime}_{n,n_{0}}$ coincides with the initial one $b_{n,n_{0}}$, because of the continuity of the implicit function theorem used to extend the curve in Proposition 13. In this way one creates a curve which is a graph of a smooth function. This is obtained by using the angle estimate in Proposition 4 and Lemma 14. Because of lemmas 15 and 16 the curve built cannot exit the interior of $\mathcal{B}_{n}$. ∎
@@
Let $U$ be a neighborhood of $b_{n,n_{0}}$ which is strictly contained in $\mathcal{B}_{n}$. Observe that the new family is a restriction of our original family and the secondary tangency curve $b_{n,n_{0}}$ describes homoclinic tangencies associated to the original saddle point. The transversal intersection $q_{2}$ it is still present in the new family. As a consequence, conditions (f1),(f2),(f3),(f4),(f5),(f6),(F1),(F2),(F3),(F5)(f1),(f2),(f3),(f4),(f5),(f6),(F1),(F2),(F3),(F5) are automatically satisfied.
Observe that $f_{n,n_{0}}^{n-n_{0}}\left(q_{1,n,n_{0}}\right)$ converges to $q_{2}$. Property $(f6)$ implies $(f7)$ for the secondary tangency $q_{1,n,n_{0}}$.
Use the notation $m_{1},m_{2},m_{3},m_{4},m_{5}$ from Proposition 2 and observe that $m_{1}\in W^{u}_{\text{loc}}(0)$ close to $q_{3}$. This proves $(f8)$ for the secondary tangency $q_{1,n,n_{0}}$.
@@
Observe that $(P1)$ follows by the fact that $b_{n,n_{0}}$ is the curve of tangencies and $(P2)$ from the fact that $\Psi_{1,3}\neq 0$, see Lemma 13, Equation (5.1). ∎
@@ 6 Newhouse phenomenon
In this section, we identify parameters corresponding to maps that exhibit infinitely many sinks, a phenomenon referred to as the Newhouse phenomenon. Although this result itself is not new, our approach is constructive in nature and departs fundamentally from earlier methods, which in turn enables us to establish a novel and unexpected property of unfoldings.
We start by proving the existence of sinks in two-dimensional unfoldings.
@@ 6.1 Cascades of sinks
In this section we are going to prove that, for a two dimensional set of parameters, an unfolding contains maps which have a sink of high period. Choose $\epsilon_{0}>0$ and define, for $n$ large enough,
@@
Remark 11.
Since $\partial\mu/\partial a$, $\partial\lambda/\partial t$ and $\partial\lambda/\partial a$ are all bounded we have that, for all $(t,\tilde{a})\in\mathcal{A}_{n}$,
$\left(\frac{\mu(t,a_{n}(t))}{\mu(t,\tilde{a})}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right),\left(\frac{\lambda(t,a_{n}(t))}{\lambda(\tilde{t},\tilde{a})|}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right).$
@@
For $n$ large enough, take $(t,a)\in\mathcal{A}_{n}$ and denote by $c_{n}(t)=c_{t,a_{n}(t)}$ and by $z_{n}(t)=z_{t,a_{n}(t)}=(z_{n,x}(t),z_{n,y}(t))$. When the choice of $t$ is clear, we just use the notation $c_{n}$ and $z_{n}$.
For $\delta>0$ we define the box that is going to contain the sink,
@@
When the choice of $(t,a)$ is clear we just use the notation $B^{n}_{\delta}$.
In the next lemma we prove that $B^{n}_{\delta}$ returns into itself, see Figure 7. Let $\tilde{Q}=\max Q_{x,y}$ where $Q_{x,y}$ as in Lemma 5.
@@
[…]
$|\Delta x|\leq\frac{1}{3},|\Delta y|\leq\frac{\delta}{\mu(t,a_{n}(t))^{2n}}\text{ and }|\Delta a|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}.$ (6.1)
Denote by $\left(\Delta x^{\prime},\Delta y^{\prime}\right)=F^{n}_{t,a}\left(\Delta x,\Delta y\right)-c_{t,a}$. Using that $F^{n}_{t,a}$ is linear on $B^{n}_{\delta}$, see (2.4), we get
$|\Delta x^{\prime}|\leq 2\lambda(t,a)^{n},$
@@
$|\Delta y^{\prime}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+\left|\mu(t,a)^{n}z_{n,y}-c(0,t,a)\right|$
$\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+\left|\left(\left[1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)\right]\mu(t,a_{n}(t))^{n}z_{n,y}\right)-c_{n,y}\right|$
$+$ $\left|c_{n,y}-c(0,t,a)\right|,$
where we used Remark 11. Since $F^{n}_{t,a_{n}(t)}\left(z_{n}\right)\in\Gamma_{t,a_{n}(t)}$, then
$\mu(t,a_{n}(t))^{n}z_{n,y}=c_{n,y}+O\left(\lambda(t,a_{n}(t))^{n}\right).$
Moreover $\left|c_{n,y}-c(0,t,a)\right|=O\left(\Delta a\right)$ (see Lemma 1). Hence, by (6.1) and by Remark 11 we get
$|\Delta y^{\prime}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y\right|+O\left(\lambda(t,a_{n}(t))^{n}\right)+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)+O\left(\frac{1}{\mu(t,a_{n}(t))^{2n}}\right)$
@@
By Lemma 5 (center $F^{N}_{t,a}$ in coordinates around $c_{t,a}$) extended to include also the Taylor expansion in $\Delta a$ we get, for $n$ large enough
[…]
@@
[…]
$\leq$ $\frac{\tilde{Q}\delta^{2}}{\mu(t,a_{n}(t))^{2n}}+O\left(\frac{1}{\mu(t,a_{n}(t))^{3n}}\right)+\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}},$
where we also used $(F3)$ and Remark 11. By our choice of $\epsilon_{0}$ and $\delta$, for $n$ large enough, the lemma follows. ∎
We fix $\epsilon_{0},$ and $\delta$ such that Lemma 18 holds. We are now ready to prove the existence of a sink. This is achieved in the next proposition by proving that $F^{n+N}$ contracts the box $B_{\delta}^{n}$.
Proposition 5.
For $n$ large enough and for all $(t,a)\in\mathcal{A}_{n}$, $B^{n}_{\delta}(t)$ has a unique periodic point which is a sink.
@@
Because $F_{t,a}^{n}$ is linear on $B_{\delta}^{n}$, the image $F_{t,a}^{n}\left(B_{\delta}^{n}\right)$ is contained in a neighborhood of $c_{t,a}$ of diameter smaller than $\delta\mu^{-n}\left(1+O\left({n}\mu^{-n}\right)\right)$. From this and by differentiating (2.15) with respect to $\Delta x$ and $\Delta y$ we get
@@
$DF_{t,a}^{n+N}$ $=$ $\left(\begin{matrix}O(1)&O(1)\\ O(1)&{2\delta}{\mu^{-n}}Q_{x,y}\left(1+O\left({n}{\mu^{-n}}\right)\right)\end{matrix}\right)\left(\begin{matrix}O\left(\lambda^{n}\right)&0\\ 0&\mu^{n}\end{matrix}\right)$
@@ 6.2 Newhouse phenomenon: infinitely many sinks
We are now ready to set up an induction procedure to prove the existence of maps having infinitely many sinks. This procedure uses the so-called “Newhouse boxes”. In the first generation, the Newhouse boxes are essentially rectangles in $\mathcal{A}_{n}$ whose boundary are defined by the curves of secondary tangencies, see Figure 6. The family restricted to the Newhouse boxes of first generation have one sink and it is an unfolding of a new homoclinic tangency. The propositions and the lemmas proved in the previous sections apply then to these families creating Newhouse boxes of second generation. As a consequence, the family restricted to the Newhouse boxes of second generation have two sinks and it is an unfolding of a new homoclinic tangency. We proceed by induction.
Let $(t_{n,n_{0}},a_{n,n_{0}})$ be the parameters at the intersection point $a_{n}\cap b_{n,n_{0}}$, see Proposition 4, and let
@@
Recall that $b_{n,n_{0}}$ is the graph of a function $b=b_{n,n_{0}}:[-t_{0},t_{0}]\mapsto\mathbb{R}$. Consider its restriction to the interval $t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}$] where $t^{-}_{n,n_{0}}$ and $t^{+}_{n,n_{0}}$ are the two intersection points of the curve $b$ with the boundary of the strip $\mathcal{A}_{n}$, see Figure 6.
The domains
$\mathcal{P}_{n,n_{0}}=\left\{(t,a)\left|\right.t\in[t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}],|a-a_{n}(t)|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}\right\}$
@@
By Proposition 4 the curve $b_{n,n_{0}}$ is the graph of the function $b$ and by the angle estimate in Lemma 13, there exists a uniform constant $K>0$ such that
$\frac{1}{K}\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}\leq\left|\left\{t\left|\right.(t,b(t))\in\mathcal{A}_{n}\right\}\right|\leq K\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}.$ (6.2)
Moreover the proof of Proposition 3 gives $\text{dist}\left(f_{n,n_{0}},f_{n,n_{0}+1}\right)$ is proportional to 1/n{1}/{n}. The disjointness follows from this estimates, the fact that $\lambda^{\theta}\mu^{2}\geq\left(\lambda^{2\theta}\mu^{3}\right)^{1/2}>1$, see (2.16) and the fact that $\mathcal{A}_{n}$ are pairwise disjoint. ∎
The next proposition ensures that the family restricted to $\mathcal{P}_{n,n_{0}}$ is again an unfolding. This allows an inductive procedure.
Proposition 6.
For $n$ large enough, the map $f_{n,n_{0}}$ has a strong homoclinic tangency and the restriction $F:\mathcal{P}_{n,n_{0}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become an unfolding. Moreover each map in $\mathcal{P}_{n,n_{0}}$ has a sink of period $n+N$.
@@
The proof is given in Lemma 17. ∎
Inductively we are going to construct parameters with multiple sinks of higher and higher periods and a strong homoclinic tangency. Let $\mathfrak{N}^{1}=\left\{(n,n_{0})\left|\right.n_{0}\in\mathfrak{N}_{n}\right\}\subset{\mathbb{N}}^{2}$ be the set of labels of the Newhouse boxes of first generation $\mathcal{P}_{n,n_{0}}$. As inductive hypothesis assume that there exist sets $\mathfrak{N}^{k}\in\left({\mathbb{N}}^{2}\right)^{k}$ such that, each $\mathfrak{N}^{k}$ is the set of labels of the Newhouse boxes of generation $k$. Moreover the natural projections $\mathfrak{N}^{1}\leftarrow\mathfrak{N}^{2}\leftarrow\dots\leftarrow\mathfrak{N}^{g}$ with $\mathfrak{N}^{k}\in\left({\mathbb{N}}^{2}\right)^{k}$ are countable to 11 and correspond to inclusion of Newhouse boxes of successive generations. They satisfy the following inductive hypothesis.
For $\underline{n}=\left\{(n^{(k)},n_{0}^{(k)})\right\}_{k=1}^{g}\in\mathfrak{N}^{g}$, there exist sets $\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}\subset\mathcal{P}$, where $n^{(k)}$ labels the sink and $n_{0}^{(k)}$ labels the secondary tangency, such that
@@
  • the restriction $F:{\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become and unfolding of $f^{k}_{n^{(k)},n_{0}^{(k)}}$,
@@
By induction, using Proposition 6, we get an infinite sequence of sets $\mathfrak{N}^{k}$.
@@
$NH=\bigcap_{g}\bigcup_{\underline{n}\in\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}.$
@@
$\overline{NH}\cap\mathcal{P}_{n,n_{0}}\supset\text{\rm graph}(b_{n,n_{0}}).$ (6.3)
@@
Given $(t,0)$, by Proposition 3, for every $n$ large enough, $a_{n}$ has a secondary tangency at $(t_{n},a_{n}(t_{n}))$ with $|t-t_{n}|=O\left(\frac{1}{n}\right)$. Hence, there exists a sequence of Newhouse boxes in $\mathfrak{N_{1}}$ accumulating at $(t,0)$. By construction, each box in $\mathfrak{N_{1}}$ contains points of $NH$. The lemma follows. ∎
@@
The inductive construction, using Proposition 6, implies that all maps in $NH$ have infinitely many sinks. From Lemma 19 we know that
$\bigcup_{\underline{n}\in\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}},$

### v3

@@ 5.3 Curves of secondary tangencies with uniform size
Let $(t_{n_{0}},a_{n_{0}})\in\mathcal{B}_{n}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. By Lemma 5.1.1 there exist $\epsilon>0$ and a smooth curve
$b_{n,n_{0}}:[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]\mapsto\mathbb{R}$
such that, for all $t\in[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$. We want now to prove that this curve does not stop. Suppose it does. Let $(t_{-},t_{+})$ be the maximal interval in $t_{0},t_{0}]$ to which this function extends smoothly such that its graph is a locus of secondary tangency of type $n_{0}$. Suppose that $t_{-},t_{+}]\subset(-t_{0},t_{0})$. Consider a sequence $t_{n}\in(t_{-},t_{+})$ of secondary tangencies of type $n_{0}$ that converges to a boundary point in $t_{-},t_{+}]$. Since the set of tangencies is a close set, all limit points are again a tangency. Because of Lemma 5.2.1 one cannot have more than one limit point, indeed one can have only one tangency point while moving in the $a$ direction. By lemmas 5.2.2 and 5.2.4 this unique limit point cannot be on the boundary lines that define $\mathcal{B}_{n}$. Such a limit point is then contained in $\mathcal{B}_{n}$. One can then apply again Proposition 5.1.1 to this point and build a smooth curve of maps with secondary tangencies of type $n_{0}$. This new courve $b^{\prime}_{n,n_{0}}$ coincides with the initial one $b_{n,n_{0}}$, because of the uniqueness assured by the implicit function theorem used to create the local curve in Proposition 5.1.1. This contradicts the maximality of the interval $(t_{-},t_{+})$. In particular the function extends on the all domain $t_{0},t_{0}]$ without exiting $\mathcal{B}_{n}$. ∎
In the following lemma we prove that the estimate for the angle between the curve $b_{n,n_{0}}$ and $a_{n}$, see Lemma 5.1.1, persists in a neighborhood of $a_{n}$ of width $O\left(1/\mu^{2n}\right)$.
Lemma 5.3.2.
For all $t\in[-t_{0},t_{0}]$, with $(b_{n,n_{0}}(t)-a_{n}(t))=O\left(1/\mu^{2n}\right)$,
$\frac{db_{n,n_{0}}}{dt}=\frac{da_{n}}{dt}+{n\lambda^{\theta n}}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta n}\right).$
Proof.
Observe that, by (2.3.1), (4.2.6) holds for all parameters $(t,a)$ with $a=a_{n}+O\left(1/\mu^{2n}\right)$. As a consequence (4.2) and (4.2.31) hold as well and this implies that (5.1.3) and (5.1.4) also hold. In particular (5.1.6) holds for all parameters $(t,a)$ with $a=a_{n}+O\left(1/\mu^{2n}\right)$. The proof is complete. ∎
@@
Let $U$ be a neighborhood of $b_{n,n_{0}}$ which is strictly contained in $\mathcal{B}_{n}$. Observe that the new family is a restriction of our original family and the secondary tangency curve $b_{n,n_{0}}$ describes homoclinic tangencies associated to the original saddle point. The transversal intersection $q_{2}$ it is still present in the new family. As a consequence, conditions (f1),(f2),(f3),(f5),(f6),(F1),(F2),(F3),(F5)(f1),(f2),(f3),(f5),(f6),(F1),(F2),(F3),(F5) are automatically satisfied. Condition $(f4)$ is a consequence of Lemma 4.2.36. By construction, condition $(f6)$ implies $(f7)$ for the secondary tangency $q_{1,n,n_{0}}$.
Use the notation from Proposition 4.2.2 and observe that $m_{1}\in W^{u}_{\text{loc}}(0)$ close to $q_{3}$ such that $F^{3N+\theta n+n+n_{0}}(m_{1})=q_{1,n,n_{0}}$. This proves $(f8)$ for the secondary tangency $q_{1,n,n_{0}}$.
@@
Observe that $(P1)$ follows by the fact that $b_{n,n_{0}}$ is the curve of non-degenerate tangencies, see Lemma 4.2.36 and $(P2)$ from Lemma 5.2.1. ∎
@@ 6 Newhouse phenomenon
In this section, we identify parameters corresponding to maps that have infinitely many sinks of arbitrarily high period, a phenomenon referred to as the Newhouse phenomenon. Although this result itself is not new, our approach is constructive in nature and departs fundamentally from earlier methods, which in turn enables us to describe the dynamics on the closure of these sinks establishing a novel and unexpected property of unfoldings. We start by proving the existence of sinks in two-dimensional unfoldings.
@@ 6.1 Cascades of sinks
In this section we are going to prove that, for a specific open set of parameters, the corresponding maps have a sink of high period. Choose $\epsilon_{0}>0$ and define, for $n$ large enough,
@@
Remark 6.1.1.
As in Remark 4.0.3, for all $(t,\tilde{a})\in\mathcal{A}_{n}$,
$\left(\frac{\mu(t,a_{n}(t))}{\mu(t,\tilde{a})}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right),\left(\frac{\lambda(t,a_{n}(t))}{\lambda(\tilde{t},\tilde{a})}\right)^{n}=1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right).$
@@
For $n$ large enough, take $(t,a)\in\mathcal{A}_{n}$ and denote by $c_{n}(t)=c_{t,a_{n}(t)}$ and by $z_{n}(t)=z_{t,a_{n}(t)}=(z_{n,x}(t),z_{n,y}(t))$. When the choice of $t$ is clear, we just use the notation $c_{n}$ and $z_{n}$. For $\delta>0$ we define the box that is going to contain the sink,
@@
When the choice of $(t,a)$ is clear we just use the notation $B^{n}_{\delta}$. In the next lemma we prove that $B^{n}_{\delta}$ returns into itself, see Figure 7. Let $\tilde{Q}=\max Q_{x,y}$ where $Q_{x,y}$ is as in Lemma 2.2.12.
@@
Fix $(t,a)\in\mathcal{A}_{n}$. For $n$ large enough, we write $F^{n+N}_{t,a}$ in coordinates centered at $z_{n}$, namely $F^{n+N}_{t,a}\left(\Delta x_{z_{n}},\Delta y_{z_{n}}\right)=\left(\Delta\tilde{x}_{z_{n}},\Delta\tilde{y}_{z_{n}}\right)$. Let $\left(\Delta x_{z_{n}},\Delta y_{z_{n}}\right)$ be such that if $z_{n}+\left(\Delta x_{z_{n}},\Delta y_{z_{n}}\right)\in B^{n}_{\delta}$ and $\Delta a=(a-a_{n}(t))$, then
$|\Delta x_{z_{n}}|\leq\frac{1}{3},|\Delta y_{z_{n}}|\leq\frac{\delta}{\mu(t,a_{n}(t))^{2n}}\text{ and }|\Delta a|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}.$ (6.1.3)
Denote by $\left(\Delta x_{c_{t,a}},\Delta y_{c_{t,a}}\right)=F^{n}_{t,a}\left(\Delta x_{z_{n}},\Delta y_{z_{n}}\right)-c_{t,a}$. Using that $F^{n}_{t,a}$ is linear on $B^{n}_{\delta}$, see (2.1.8), we get
$|\Delta x_{c_{t,a}}|\leq 2\lambda(t,a)^{n},$
@@
$|\Delta y_{c_{t,a}}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y_{z_{n}}\right|+\left|\mu(t,a)^{n}z_{n,y}-c(0,t,a)\right|$
$\leq$ $\mu(t,a)^{n}\left|\Delta y_{z_{n}}\right|+\left|\left(\left[1+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)\right]\mu(t,a_{n}(t))^{n}z_{n,y}\right)-1\right|,$
where we used Remark 6.1.1. Since $F^{n}_{t,a_{n}(t)}\left(z_{n}\right)\in\Gamma_{t,a_{n}(t)}$, then
$\mu(t,a_{n}(t))^{n}z_{n,y}=1+O\left(\lambda(t,a_{n}(t))^{n}\right).$
Hence, by (6.1.3) and by Remark 6.1.1 we get,
$|\Delta y_{c_{t,a}}|$ $\leq$ $\mu(t,a)^{n}\left|\Delta y_{z_{n}}\right|+O\left(\lambda(t,a_{n}(t))^{n}\right)+O\left(\frac{n}{\mu(t,a_{n}(t))^{2n}}\right)$
@@
By Lemma 2.2.12 (center $F^{N}_{t,a}$ in coordinates around $c_{t,a}$) extended to include also the Taylor expansion in $\Delta a$ we get, for $n$ large enough
$|\Delta\tilde{x}_{z_{n}}|$ $=$ $O\left(\Delta x_{c_{t,a}}\right)+O\left(\Delta y_{c_{t,a}}\right)+O\left(\Delta a\right)\leq\frac{1}{3},$
@@
$|\Delta\tilde{y}_{z_{n}}|$ $\leq$ $O\left(\lambda(t,a)^{n}\right)+Q_{x,y}\left|\Delta y_{c_{t,a}}\right|^{2}+O\left(\frac{1}{\mu(t,a_{n}(t))^{3n}}\right)+|\Delta a|$
$\leq$ $\frac{\tilde{Q}\delta^{2}}{\mu(t,a_{n}(t))^{2n}}+O\left(\frac{n}{\mu(t,a_{n}(t))^{3n}}\right)+\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}},$
where we also used $(F3)$ and Remark 6.1.1. By our choice of $\epsilon_{0}$ and $\delta$, for $n$ large enough, the lemma follows. ∎
We fix $\epsilon_{0},$ and $\delta$ such that Lemma 6.1.2 holds. We are now ready to prove the existence of a sink. This is achieved in the next proposition by proving that $F^{n+N}$ contracts the box $B_{\delta}^{n}$.
Proposition 6.1.4.
For $n$ large enough and for all $(t,a)\in\mathcal{A}_{n}$, $B^{n}_{\delta}(t)$ has a unique periodic point which is a sink of period $n+N$.
@@
Because $F_{t,a}^{n}$ is linear on $B_{\delta}^{n}$, the image $F_{t,a}^{n}\left(B_{\delta}^{n}\right)$ is contained in a neighborhood of $c_{t,a}$ of diameter smaller than $\delta\mu^{-n}\left(1+O\left({n}\mu^{-n}\right)\right)$. From this and by differentiating (2.2.13) with respect to $\Delta x$ and $\Delta y$ we get
@@
$DF_{t,a}^{n+N}$ $=$ $\left(\begin{matrix}O(1)&O(1)\\ O(1)&{2\delta}{\mu^{-n}}Q_{x,y}\left(1+O\left({n}{\mu^{-n}}\right)\right)\end{matrix}\right)\left(\begin{matrix}\lambda^{n}&0\\ 0&\mu^{n}\end{matrix}\right)$
@@ 6.2 Newhouse phenomenon: infinitely many sinks
We are now ready to set up an induction procedure to prove the existence of maps having infinitely many sinks. This procedure uses “Newhouse boxes”. In the first generation, the Newhouse boxes are essentially rectangles in $\mathcal{A}_{n}$ whose boundary are defined by the curves of secondary tangencies, see Figure 6. The family restricted to the Newhouse boxes of first generation have one sink and it is an unfolding of a new homoclinic tangency. The propositions and the lemmas proved in the previous sections apply then to these families creating Newhouse boxes of second generation. As a consequence, the family restricted to the Newhouse boxes of second generation have two sinks and it is an unfolding of a new homoclinic tangency. We proceed by induction.
Let $(t_{n,n_{0}},a_{n,n_{0}})$ be the parameters at the intersection point $a_{n}\cap b_{n,n_{0}}$, see Proposition 5.3.1, and let
@@
Recall that $b_{n,n_{0}}$ is the graph of a function $b=b_{n,n_{0}}:[-t_{0},t_{0}]\mapsto\mathbb{R}$. Consider its restriction to the interval $t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}$] where $t^{-}_{n,n_{0}}$ and $t^{+}_{n,n_{0}}$ are the two intersection points of the curve $b$ with the boundary of the strip $\mathcal{A}_{n}$, see Figure 6. The existence of these intersection points is guaranteed by Lemma 5.3.2. The domains,
$\mathcal{P}_{n,n_{0}}=\left\{(t,a)\left|\right.t\in[t^{-}_{n,n_{0}},t^{+}_{n,n_{0}}],|a-a_{n}(t)|\leq\frac{\epsilon_{0}}{\mu(t,a_{n}(t))^{2n}}\right\},$
@@
By Proposition 5.3.1 the curve $b_{n,n_{0}}$ is the graph of the function $b$ and by the angle estimate in Lemma 5.3.2, there exists a uniform constant $K>0$ such that
$\frac{1}{K}\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}\leq\left|\left\{t\left|\right.(t,b(t))\in\mathcal{A}_{n}\right\}\right|\leq K\frac{1}{n\left(\lambda^{\theta}\mu^{2}\right)^{n}}.$ (6.2.2)
Moreover the proof of Proposition 4.2.32 gives $\text{dist}\left(f_{n,n_{0}},f_{n,n_{0}+1}\right)$ is proportional to 1/n{1}/{n}. The disjointness follows from this estimates, the fact that $\lambda^{\theta}\mu^{2}\geq\left(\lambda^{2\theta}\mu^{3}\right)^{1/2}>1$, see (2.3.1) and the fact that $\mathcal{A}_{n}$ are pairwise disjoint. ∎
The next proposition ensures that the family restricted to $\mathcal{P}_{n,n_{0}}$ is again an unfoldings. This allows an inductive procedure.
Proposition 6.2.3.
For $n$ large enough, the map $f_{n,n_{0}}$ has a strong homoclinic tangency and the restriction $F:\mathcal{P}_{n,n_{0}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become an unfoldings. Moreover each map in $\mathcal{P}_{n,n_{0}}$ has a sink of period $n+N$.
@@
The proof is given in Lemma 5.3.3 and Proposition 6.1.4. ∎
Inductively we are going to construct parameters with multiple sinks of higher and higher periods and a strong homoclinic tangency. Let $\mathfrak{N}^{1}$ be the set of labels of the Newhouse boxes of first generation, i.e. all possible pairs $(n,n_{0})$ for which there exists a Newhouse box $\mathcal{P}_{n,n_{0}}$. As inductive hypothesis assume that there exist sets $\mathfrak{N}^{k}\subset{\mathbb{N}}^{2}$ such that, each $\mathfrak{N}^{k}$ is the set of labels of the Newhouse boxes of generation $k$. Moreover the natural projections $\mathfrak{N}^{1}\leftarrow\mathfrak{N}^{2}\leftarrow\dots\leftarrow\mathfrak{N}^{g}$ with $\mathfrak{N}^{k}\in{\mathbb{N}}^{2}$ are countable to $1$ and correspond to inclusion of Newhouse boxes of successive generations. They satisfy the following inductive hypothesis.
For $\underline{n}=\left\{(n^{(k)},n_{0}^{(k)})\right\}_{k=1}^{g}$, there exist sets $\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}\subset\mathcal{P}$, where $n^{(k)}$ labels the sink and $n_{0}^{(k)}$ labels the secondary tangency, such that
@@
  • the restriction $F:{\mathcal{P}^{k}_{n^{(k)},n_{0}^{(k)}}}\times\mathcal{M}\to\mathcal{M}$ can be reparametrized to become and unfoldings of $f^{k}_{n^{(k)},n_{0}^{(k)}}$,
@@
By induction, using Proposition 6.2.3, we get an infinite sequence of sets $\mathfrak{N}^{k}$.
@@
$NH=\bigcap_{g}\bigcup_{\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}}.$
@@
$\overline{NH}\cap\mathcal{P}_{n,n_{0}}\supset\text{\rm graph}(b_{n,n_{0}})\cap\mathcal{P}_{n,n_{0}}.$ (6.2.6)
@@
Given $(t,0)$, by Proposition 4.2.32, for every $n$ large enough, $a_{n}$ has a secondary tangency at $(t_{n},a_{n}(t_{n}))$ with $|t-t_{n}|=O\left(\frac{1}{n}\right)$. Hence, there exists a sequence of Newhouse boxes in $\mathfrak{N_{1}}$ accumulating at $(t,0)$. By construction, each box in $\mathfrak{N_{1}}$ contains points of $NH$. The lemma follows. ∎
@@
The inductive construction, using Proposition 6.2.3, implies that all maps in $NH$ have infinitely many sinks. From Lemma 6.2.1 we know that
$\bigcup_{\mathfrak{N}^{g}}\mathcal{P}^{g}_{n^{(g)},n_{0}^{(g)}},$